• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 00:16:11 on February25,2009

    Entry number 263196

    keyword=Swing shift summary

    Swing shift summary:

    shift workers: J. P. Chen (SL), D. Wang (TO), X. Deng (3rd)

    16:00 Beam off due to RF problem
    16:45 beam back, started with 5uA, checked rates, then 15 uA.
    Continue production for kin 4K.
    17:50 BCM ioc (IOCHA5) not working. Rebooted (from the middle room).
    17:55 Put back correct BCm calibration constant.
    18:40, Compton HV tripped. Reset.
    22:03 change LHRS from 1.6 to 1.51 GeV/c (kin. 4J), starting from run
    20390/1777.
    22:10 moved targte to reference cell: N2 113 psig.
    22:45 moved to pol. He3.

    LHRS runs:
    20386 16:49-18:34, 400K
    20387, 18:35-20:13, 400K
    20388, 20:13-21:50, 400K
    20389, 21:50-22:02, 43K

    Kin 4J.
    20390, short run on pol. He3
    20391, 22:14-22:45, 117K, N2
    20392, 22:50- still ongoing at 24:00. He3.

    BB runs:
    1773, 16:49-18:21, 7M
    1774, 18:22-19:32, 7M
    1775, 19:33-20:48, 7M
    1776, 20:48-22:02, 7M

    1777, short run on pol. He3
    1778, 22:14-22:45, 2.8M, N2.
    1779, 22:50-0:03,7M, He3.